The sum of the ages of a father and son is 56. Four years ago the father was 3 times as old as the son. Find the present age of each.
Answer by macston(5194)

F=father; S=son
F+S=56yrs
F=56yrs-S
F-4yrs=3(S-4yrs) Substitute for F.
56yrs-S-4yrs=3S-12 yrs Add S+12yrs to each side.
64yrs=4S Divide each side by 4.
16yrs=S ANSWER 1: The son is 16 years old.
F=56yrs-S=56yrs-16yrs=40yrs ANSWER 2: The father is 40 years old.
CHECK:
F-4yrs=3(S-4yrs)
40yrs-4yrs=3(16yrs-4yrs)
36yrs=36yrs